# ConRebSeg: A Segmentation Dataset for Reinforced Concrete Construction

### v1

@@ 3.5 Annotation Styles / Cleanup
A phenomenon we observed while analyzing the segmentation masks is that the same object has been annotated in different ways. This was to be expected since we shuffled the frames within sequences collected at Langebro (see Table 2) to remove the temporal dependencies within frames since we wanted to analyze how different annotators will label the ExposedBars category. Figure 5 shows three consequent frames with minimal changes depicting the same object, along with the mask.
Figure 5: Sequence of frames with its labels showing inconsistency in labeling style
It can be seen that the mask drastically changes in appearance, moving from a continuous ”area” annotation to a more fine-grained, ”rebar” annotation. While our label description explained that ”areas” should be annotated, we conclude that both styles are valid annotations that serve different purposes. The styles can be considered ”expert” opinions, and both need to be respected. However, we assume that two vastly different masks for very similar images might impede the training of deep learning-based segmentation models. We will elaborate on this in Section 4.4 in more detail, and describe an approach to separate the different styles in the following.
Figure 5 suggests an indicator being the number of objects in an image. The object count from the left picture in Figure 5 to the middle one suddenly increases from 1 to 7 and then goes back to 1 in the last picture. We, therefore, extract the number of unconnected masks per frame and create a time series from it. A local inconsistency is then determined by a rolling window approach of taking $n=10$ frames centered around the current frame, determining their 0.90 quantiles $q_{.90}$ of the object count, and marking a frame as an anomaly if the number of objects in the frame exceeds $q_{.90}$. Formally, the rule is expressed as follows:
$\text{anomaly}=\begin{cases}\text{1},&\text{if }n_{\text{obj}}>q_{.90}\\ \text{0},&\text{otherwise}\end{cases}$ (1)
with 1 indicating an anomaly and 0 indicating that something is adhering to the ”common” style. We now use the train sequence run_2023_06_07_10_05_31 from the Langebro sequences as an example. Figure 6 shows a plot of the time series, with the number of objects in a frame noted on the y-axis. The green dashed line indicates $q_{.90}$, and any point above this line is considered an anomaly and is denoted as an orange dot. Based on this rule, we tag the frames within a sequence with an anomaly tag if they fulfill it.
Figure 6: Plot of the number of objects per frame in the Langebro sequence run_2023_06_07_10_05_31. Orange dots indicate anomalies. The dashed green line indicates the .90 quantile $q_{.90}$
Figures 7(a) and 7(b) show the result of the tagging process on Langebro sequence
@@
Figure 7: Comparison between frames without the anomaly tag (left) and with the anomaly tag (right)
@@ 4 Experiments
This section presents the various experiments conducted on the dataset. We will first present our infrastructure used in the experiments. Then, we will establish some baselines for different semantic and instance segmentation models, and carry out experiments regarding data efficiency and the influence of certain label styles on the performance of the models.
@@ 4.1 Setup
For our experiments, we use three different setups as listed in Table 4. All of the models used in the experiments are based on PyTorch [48] and trained on CUDA-enabled GPUs. To ensure consistent and comparable results within and across models, we enable deterministic training and initialize the random number generators with fixed seeds.
@@ 4.2 Baseline Training
In the first round, we train from scratch with default optimizer hyperparameters, and we only consider data from the first 100 epochs. For all models, we use an image size of 640x640 pixels. YOLOv8L-seg uses extensive image augmentations, whereas DeepLabV3 and U-Net use only RandomHorizontalFlip and RandomVerticalFlip augmentations, both with $p=0.5$.
@@
Table 5 shows the results of the baseline experiments. While DeepLabV3 and U-Net achieve similar performances with regards to mean Intersection Over Union (mIOU) on the val split (mIOU $\approx$ 0.3), YOLOv8L-seg almost performs a magnitude better with a val mIOU of approximately 0.5. Given the superior performance of YOLOv8L-seg compared to the other two models, we proceed to conduct further experiments based on the YOLOv8L-seg model. A straightforward experiment is to start the training from pre-trained weights provided by the developers of the model. The result of this experiment is listed in Table 6 and shows improvement compared to the model trained from scratch. Therefore, trainings we conduct in further experiments will also be started from pre-trained weights.
@@ 4.3 Effect of withheld training data
With the established baselines being trained on 100% of the training data, we want to analyze the influence of withheld data on the performance of the models. This gives an intuition about how much variation the data introduces and how diverse our dataset is. There are two ways to withhold the data; one can either withhold whole training sequences, or withhold samples from every training sequence. We opted for the former to actually reduce the variation in data.
@@
Figure 8: Instance class distribution, in absolute values (left) and relative values (right)
Figure 9: Best mask mAP50-95 value achieved during training, measured on the val split, per portion of withheld training data
Figure 10: mIOU value of models trained on a percentage of withheld training data, on the full train split (left) and val split (right)
@@
4.4 Effect of different expert opinions
Starting from the tagging process for anomalous samples as outlined in Section 3.5, we analyze the influence of the presence of different annotation styles associated with visually similar samples. We withhold Langebro training samples tagged as anomalies followed by a training of the YOLOv8L-seg model with the same training settings as the established baselines. To ensure a fair comparison, we also consider the amount of withheld data and use the results from Section 4.3 to have a second, more appropriate baseline to compare against. In total, by applying the rule outlined in Section 3.5, we omit 150 samples from the train split. The effects on the instance counts are outlined in Table 7. The results of this experiment are listed in Table 8.
@@ 4.4 Effect of different expert opinions
We verify this by conducting another experiment, constructing a batch of seven consecutive samples (frame-label pairs). From the frames, we create a mean image to have a common input. We then pass this input through a DeepLabV3 model with a ResNet50 backbone, and use pre-trained weights (ImageNet) to create outputs. We then calculate the cross-entropy loss and evaluate the gradient of each loss with respect to all of the model’s parameters and investigate the cosine similarity of the respective gradients. Conflicting samples would then show a strong negative cosine similarity, whereas supporting samples would exhibit a strong positive cosine similarity. Figure 11 shows the data used for this experiment. Figure 12 shows the mean cosine similarity between the gradients for each sample loss.
@@
Figure 11: Input data used to analyze gradient interactions between samples
Figure 12: Cosine similarities between sample loss gradients, left mean cosine similarity, right standard deviation
All values are positive, meaning that there is no gradient conflict among samples’ loss gradients. There are also no values that strike out from the general picture. Therefore, this experiment supports the previously drawn conclusion that the influence of these different label styles is minimal and can be disregarded.
@@ 5 Results
In this section, we report results for the YOLOv8L-seg model trained from pre-trained weights on 100% of the training data using default hyperparameters, as it was the best-performing model in our experiments conducted in Section 4. For this final evaluation, we are also using the test split of the dataset. We will also identify common modes of failure and present ways to mitigate these. To verify the results, we have conducted additional verification runs with different seeds and report the results (min, mean, max) in Table 9.
@@
For all metrics, the performance is roughly stable within a split with only small fluctuations. When measuring the same metrics on the different splits of the dataset, the performance naturally decreases. This is a sign that the model is overfitting and is usually alleviated by introducing more variation in the training data, accompanied by regularization techniques. Table 9 also shows that the drop from val to test is not as drastic as from train to val, indicating that the validation split is representative of the test split as well.
When analyzing the mean mask precision and recall, we see that the model performance is heavily impacted by a significant drop in recall. While the drop in precision is about half a magnitude when going from the train split to the test split, it is roughly two-thirds for the recall. This means that the model often fails to localize the objects in general. The root cause of this is usually an insufficient number of instances in the dataset. Since Table 9 shows macro metrics, it is imperative to recall that our dataset is imbalanced and does not have a lot of instances for classes other than the ExposedBars category. Thus, it makes sense to separately calculate these metrics solely for this category, which is listed in Table 10. We can see that the ExposedBars category performs better than the macro average in terms of mIOU – the test mIOU shows an improvement of almost a magnitude. The mAP50-95 on the test split is slightly higher than the macro average, probably due to the increased precision for the ExposedBars category. Compared to the macro average mean value, the mean mask precision is almost double as high when solely considering the ExposedBars category. The mean mask recall is roughly on par with the macro average mean value.
@@
  • Classification errors, i.e. properly localized bounding box, but wrong class label
@@
  • Background errors: A false positive detection, i.e. a predicted bounding box has an IOU smaller than the background threshold with a ground truth box
@@
    Figure 13: Error analysis of bounding box predictions
Figure 13 shows the results of the bounding box error analysis conducted with the TIDE tool. Throughout the splits, one big source of errors are missed detections, i.e., the model fails to produce a prediction for some ground truth instance. After this type, false positive detections are also a significant source of errors, i.e. the model recognizes unimportant backgrounds as an object. Localization and classification errors are, compared to the other error types, minimal. We can also see that $\Delta$mAP value for missed detections is almost double compared to the train split, which is in line with the mask mean recall values reported in Tables 9 and 10.

### v2

@@ 3.5 Annotation Styles / Cleanup
A phenomenon we observed while analyzing the segmentation masks is that the same object has been annotated in different ways. This was to be expected since we shuffled the frames within scenes collected at Langebro (see Table 2) to remove the temporal dependencies within frames since we wanted to analyze how different annotators will label the ExposedBars category. Figure 5 shows three consequent frames with minimal changes depicting the same object, along with the mask.
Figure 5: Sequence of successive frames with their labels (in red) showing an inconsistency in labeling style. The left and right frame share the same annotation style, the center frame has a different style.
It can be seen that the mask drastically changes in appearance, moving from a continuous ”area” annotation to a more fine-grained, ”rebar” annotation. While our label description explained that ”areas” should be annotated, we conclude that both styles are valid annotations that serve different purposes. The styles could reflect ”expert” opinions, which both need to be respected. However, we assume that two vastly different masks for very similar images might impede the training of deep learning-based segmentation models. We will elaborate on this in Section 4.4 in more detail, and describe an approach to separate the different styles in the following.
Figure 5 suggests an indicator being the temporal development of the number of objects (per-class) in an image. The object count from the left picture in Figure 5 to the middle one suddenly increases from 1 to 7 and then goes back to 1 in the last picture. We, therefore, extract the number of unconnected masks of the same class per frame and create a time series from it, respecting the temporal order of the frames. A local inconsistency is then determined by a rolling window approach of taking $n=10$ frames centered around the current frame, determining their 0.90 quantiles $q_{.90}(t)$ of the object count, and marking a frame as an anomaly if the number of objects in the frame exceeds $q_{.90}(t)$. Formally, the rule is expressed as follows:
$\text{anomaly}(t)=\begin{cases}\text{1},&\text{if }n_{\text{obj}}(t)>q_{.90}(t)\\ \text{0},&\text{otherwise}\end{cases}$ (1)
with 1 indicating an anomaly and 0 indicating that something is adhering to the ”common” style. We now use the train scene run_2023_06_07_10_05_31 from the Langebro scenes as an example. Figure 6 shows a plot of the time series, with the number of objects $n_{obj}(t)$ in a frame noted on the y-axis. The green dashed line indicates $q_{.90}(t)$, and any point above this line is considered an anomaly and is denoted as an orange dot. Based on this rule, we tag the frames within a scene with an anomaly tag if they fulfill it.
Figure 6: Plot of the number of objects $n_{obj}(t)$ (on y-axis) per frame (on x-axis) in the Langebro scene run_2023_06_07_10_05_31. Orange dots indicate anomalies. The dashed green line indicates the local .90 quantile $q_{.90}(t)$
Figures 7(a) and 7(b) show the result of the tagging process on Langebro scene
@@
Figure 7: Comparison between frames and their labels (in yellow) without the anomaly tag (left) and with the anomaly tag (right)
@@ 4 Experiments
This section presents the various experiments conducted on the dataset. We will first present our infrastructure used in the experiments. Then, we will establish some baselines for different semantic and instance segmentation models, and carry out experiments regarding data efficiency and the influence of certain labelling styles on the performance of the models.
@@ 4.1 Setup
For our experiments, we use three different setups, as listed in Table 4. All the models used in the experiments are based on PyTorch [53] and trained on CUDA-enabled GPUs. To ensure consistent and comparable results within and across models, we enable deterministic training and initialize the random number generators with fixed seeds.
@@ 4.2 Baseline Training
In the first round, we train from scratch with default optimizer hyperparameters, and we only consider data from the first 100 epochs. For all models, we use an image size of 640$\times$640 pixels. YOLOv8L-seg uses extensive image augmentations, whereas DeepLabV3 and U-Net use only RandomHorizontalFlip (RHF) and RandomVerticalFlip (RVF) augmentations, both with $p=0.5$. Table 5 summarizes the hyperparameters used for each model.
[htbp]
Table 5: Training hyperparameters used in establishing the baselines
DeepLabV3 U-Net YOLOv8L-seg
Image size 640x640 640x640 640x640
Batch size 8 8 16
Augmentations RHF and RVF with $p=0.5$ RHF and RVF with $p=0.5$ recommended default 1
Weight Initialization scratch scratch scratch
Optimizer Adam with PolynomialLR scheduler Adam SGD with momentum
Optimizer learning rate 0.001 0.001 0.01, momentum 0.9
Epochs 200 200 200
@@
Table 6 shows the results of the baseline experiments. While DeepLabV3 and U-Net achieve similar performances in regard to mean Intersection Over Union (mIOU) on the val split (mIOU $\approx$ 0.3), YOLOv8L-seg almost performs a magnitude better with a val mIOU of approximately 0.5. Given the superior performance of YOLOv8L-seg compared to the other two models, we proceed to conduct further experiments based on the YOLOv8L-seg model. A straightforward experiment is to start the training from pre-trained weights provided by the developers of the model. Those pre-trained weights stem from pre-training on the large-scale MS COCO dataset, which consists of 330,000 images [49]. The result of this experiment is listed in Table 7 and shows improvement compared to the model trained from scratch. Therefore, trainings we conduct in further experiments will also be started from pre-trained weights.
@@ 4.3 Effect of withheld training data
With the established baselines being trained on 100% of the training data, we want to analyze the influence of withheld data on the performance of the models. This gives an intuition about how much variation the data introduces and how diverse our dataset is. There are two ways to withhold the data; one can either withhold whole training scenes, or withhold samples from every training scene. We opted for the former to actually reduce the variation in data.
@@
Figure 8: Instance class distribution of train split, in absolute values (left) and relative values (right) for different amounts of withheld training data
Figure 9: Best mask mAP50-95 value (y-axis) achieved during training, measured on the val split, per portion of withheld training data (x-axis)
Figure 10: mIOU value (y-axis) of models trained on a percentage of withheld training data (x-axis), on the full train split (left) and val split (right). The purple line shows the micro average (class-agnostic mIOU), the other lines per-class mIOU.
@@
4.4 Effect of different labelling styles
Starting from the tagging process for anomalous samples as outlined in Section 3.5, we analyze the influence of the presence of different annotation styles associated with visually similar samples. We withhold Langebro training samples tagged as anomalies, followed by a training of the YOLOv8L-seg model with the same training settings as the established baselines. To ensure a fair comparison, we also consider the amount of withheld data in the train and val split and use the results from Section 4.3 to have a second, more appropriate baseline to compare against. In general, when analyzing the effect of different labelling styles on model performance, we propose the following protocol for fair comparison:
  1. 1.
    Perform experiments withholding certain amounts of training data while ensuring similar class distributions across the original and the smaller datasets.
  2. 2.
    Identify deviating labelling styles in training and validation data and remove frames containing these.
    1. (a)
      Possibly, remove more samples to ensure a similar class distribution compared to the original training and validation data
  3. 3.
    Re-train segmentation model on modified training data
  4. 4.
    Evaluate model on modified validation data
  5. 5.
    Compare obtained metrics to an appropriate baseline. Here, select the metrics of a model from Step 1 where the amount of training data is roughly the same as the amount of training data used in Step 3.
In total, applying the rule outlined in Section 3.5 and following the previously mentioned protocol, we omit 150 samples from the train and 27 samples from the val split respectively. The effect of the instance counts on the train split are outlined in Table 8, for the val split in Table 9. Relatively speaking, the different annotation styles are more explained in the train split than in the val split, however, the differences are not too big, making it fair to conclude the general effect of this phenomenon in our dataset. The results of this experiment are listed in Table 10.
@@ 4.4 Effect of different labelling styles
Table 9: Instance counts of baseline val split and with anomalous samples removed
ExposedBars person truck car       Sum
baseline 8,660 3,225 243 91 12,219
non-anomaly 8,415 3,133 243 91 11,882
Difference to baseline -245 -92 0 0 -337
relative -2.8 % -2.9 % 0 0 -2.8 %
@@
We verify this by conducting another experiment, constructing a batch of seven consecutive samples (frame-label pairs). From the frames, we create a mean image to have a common input. We then pass this input through a DeepLabV3 model with a ResNet50 backbone, and use pre-trained weights (ImageNet) to create outputs. We then calculate the cross-entropy loss and evaluate the gradient of each loss with respect to all the model’s parameters and investigate the cosine similarity of the respective gradients. Conflicting samples would then show a strong negative cosine similarity, whereas supporting samples would exhibit a strong positive cosine similarity. Figure 11 shows the data used for this experiment. Figure 12 shows the mean cosine similarity between the gradients for each sample loss.
@@
Figure 11: Input data (labels in red) used to analyze gradient interactions between samples
Figure 12: Cosine similarities between sample loss gradients, left mean cosine similarity, right standard deviation of cosine similarity. The tick labels are sample indices as listed in Figure 11
All values are positive, meaning that there is no gradient conflict among samples’ loss gradients. There are also no values that strike out from the general picture. Therefore, this experiment supports the previously drawn conclusion that the influence of these different labelling styles is minimal and can be disregarded.
@@ 5 Results
In this section, we report results for the YOLOv8L-seg model trained from pre-trained weights on 100% of the training data using default hyperparameters, as it was the best-performing model in our experiments conducted in Section 4. For this final evaluation, we are also using the test split of the dataset. We will also identify common modes of failure and present ways to mitigate these. To verify the results, we have conducted additional verification runs with different seeds and report the results (min, mean, max) in Table 11. Additionally, to reflect the robotic nature of the task, we have exported the best performing model of the verification runs to a TensorRT engine and tested the average FPS on an NVIDIA Jetson Xavier NX embedded GPU, with results reported in Table 12, as high performance desktop GPUs are often not feasible on robotic platforms.
@@
Table 12: Results of the best performing model of Table 11, when converted to TensorRT engines of different weight types (integer, half precision and full precision). FPS are reported on Jetson Xavier NX in 20W hexacore mode
INT8 FP16 FP32
train val test train val test train val test
mIOU 0.507 0.259 0.146 0.723 0.571 0.322 0.724 0.571 0.322
$\Delta\%$27.8 -56.1 -60.2 3.0 -3.2 -12.3 3.1 -3.2 -12.3
Mask mAP50-95 0.367 0.264 0.121 0.518 0.296 0.171 0.518 0.297 0.171
$\Delta\%$31.0 -11.4 -34.2 -2.6 -0.7 -7.1 -2.6 -0.3 -7.1
Mask mean precision 0.729 0.701 0.434 0.756 0.653 0.374 0.754 0.653 0.375
$\Delta\%$9.2 1.3 11.3 -5.9 -5.6 -4.1 -6.1 -5.6 -3.8
Mask mean recall 0.337 0.068 0.070 0.637 0.327 0.208 0.64 0.327 0.207
$\Delta\%$48.4 -78.9 -68.0 -2.5 1.6 -5.0 -2.0 1.6 -5.5
Avg. FPS 20.92 12.79 4.61
For all metrics, the performance is roughly stable within a split, with only small fluctuations. When measuring the same metrics on the different splits of the dataset, the performance naturally decreases. This is a sign that the model is overfitting and is usually alleviated by introducing more variation in the training data, accompanied by regularization techniques. Table 11 also shows that the drop from val to test is not as drastic as from train to val, indicating that the validation split is representative of the test split as well.
When analyzing mean mask precision and recall, we see that the model performance is heavily impacted by a significant drop in recall. While the drop in precision is about half a magnitude when going from the train split to the test split, it is roughly two-thirds for the recall. This means that the model often fails to localize the objects in general. The root cause of this is usually an insufficient number of instances in the dataset. Since Table 11 shows macro metrics, it is imperative to recall that our dataset is imbalanced and does not have a lot of instances for classes other than the ExposedBars category. Thus, it makes sense to separately calculate these metrics solely for this category, which is listed in Table 13. We can see that the ExposedBars category performs better than the macro average in terms of mIOU – the test mIOU shows an improvement of almost a magnitude. The mAP50-95 on the test split is slightly higher than the macro average, probably due to the increased precision for the ExposedBars category. Compared to the macro average mean value, the mean mask precision is almost double as high when solely considering the ExposedBars category. The mean mask recall is roughly on par with the macro average mean value. When converting the model to a TensorRT engine for deployment on an embedded GPU such as the NVIDIA Jetson Xavier NX in our case, performance drops are to be expected and also observed, as indicated in Table 12. The difference in segmentation metrics between the full floating-point precision (FP32) and half floating-point precision (FP16) are minimal, however, there are significant performance drops when using integer quantization (INT8) of the model weights. The performance drops of FP16 and FP32 are minimal, with a relative decrease of 3 % in validation mIOU, while achieving a frame rate of 12.8 and 4.6 FPS respectively. We therefore recommend using the FP16 engine as a good compromise between model accuracy and frame rate.
@@
  • 1.
    Classification errors, i.e., properly localized bounding box, but wrong class label
  • 2.
@@
  • 4.
    Background errors: A false positive detection, i.e., a predicted bounding box has an IOU smaller than the background threshold with a ground truth box
  • 5.
@@
    Figure 13: Error analysis of bounding box predictions, per split. Output from TIDE tool
Figure 13 shows the results of the bounding box error analysis conducted with the TIDE tool. Throughout the splits, one big source of errors are missed detections, i.e., the model fails to produce a prediction for some ground truth instance. After this type, false positive detections are also a significant source of errors, i.e., the model recognizes unimportant backgrounds as an object. Localization and classification errors are, compared to the other error types, minimal. We can also see that $\Delta$mAP value for missed detections is almost double compared to the train split, which is in line with the mask mean recall values reported in Tables 11 and 13.
@@
6 Discussion
Our study was aimed at introducing a new dataset for reinforced concrete construction, with a focus on public availability, contrary to most other similar works. Furthermore, it was desired to create a dataset with spatiotemporal relationships, reflecting the realistic robotic application of this dataset. The following summarizes our key findings, which will be elaborated later in this section:
  • 1.
    We have found that the performance of the segmentation model YOLOv8L-seg decreases with more and more withheld training data.
  • 2.
    We have showed a positive influence of pre-training on model performance, even when trained on unrelated, general data like the MS COCO dataset.
  • 3.
    We have showcased labelling inconsistencies in our dataset, and found that removing these inconsistencies doesn’t alter model performance significantly.
  • 4.
    We have also showed that our data has a certain center bias, i.e., the mask centroids are concentrated in the middle of the image.
  • 5.
    Our data shows high discrepancies in model performance when evaluated on a per-class level, with the car class showing low performance in terms of mIOU.
  • 6.
    An error analysis of our model revealed that false negative errors are by far the biggest error mode.
  • 7.
    On an embedded GPU, using half-floating point precision engines yields a good compromise between model accuracy and frame rate.
Our experiment showing the decrease in model performance with more and more withheld training data suggests also the opposite, i.e. model performance increases with more and more data availability. Given the lack of publicly available datasets for construction applications, we identify this property as one of the main factors why training well-performing deep learning-based models for construction applications is challenging. We therefore urge future work to make data publicly accessible to alleviate this barrier.
Because our data has a center bias in its annotations, the superiority of the YOLOv8L-seg over DeepLabV3 and U-Net can easily be explained, since YOLOv8L-seg uses a sophisticated set of image augmentations, making the model more robust to these peculiarities. Furthermore, our dataset is biased towards the ExposedBars class, with a very minimal amount of instances of the car class. This explains the discrepancies in per-class performance. This limitation can be overcome by including data from other datasets that include these classes.
The false negative errors are a clear limitation of our model. We have further analyzed this phenomenon by investigating the development of mIOU values over time in a scene and could see that some false negatives are just model instabilities, i.e. the model fails to produce a prediction for an object previously detected. Figure 14 illustrates this. This problem has already been discussed and analyzed in [9]. We stipulate that these errors are due to minor perturbations of the input resulting in big distances in feature space, thus failing to produce predictions. We therefore recommend the usage of mask stabilization techniques, as outlined in [9]. This also supports the approach of having a spatiotemporal dataset, since models are applied in real life, where one can leverage these dependencies to increase system performance.
(a) Reference frame with existing prediction, referring to the red circle in the upper plot
(b) Frame with missing prediction of same object, referring to the cyan circle in the upper plot
Figure 14: Illustration of model instabilities. Model instabilities often include a significant drop in mIOU, followed by a sudden increase in mIOU. Predictions are in cyan, ground truth annotations in orange
While our research showed minimal influences of different labelling styles on label performance, we acknowledge that this insight cannot be generalized given the low number of samples in our dataset showing this phenomenon. However, given the positive gradient interaction presented in Section 4.4, we recommend analyzing this phenomenon in other settings to verify our findings, if applicable. Furthermore, we stipulate that deep learning-based segmentation models in construction are a good use case for probabilistic segmentation models, such as Probabilistic U-Net [55], as those try to model different, valid expert opinions on how to segment a certain object. Further research should investigate the application of such models in a construction context.